OpenStudy (anonymous):

Estimate the number (0.8)^7 by using a tangent line approximation to the function f (x) = x^7 for x near the focus value a = 1. How does this estimate compare with the value given by a calculator? Compute the percentage error.

OpenStudy (amistre64):

f(1) = 1 f'(1) = 7(1)^6 = 7 tan = f'(1) (x-1) + f(1)

OpenStudy (amistre64):

at least thats how im reading this

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

bahhh i gotta eat my sammich D: Listen to amistre!! lol

OpenStudy (anonymous):

okay. I'm not sure what to do next.

OpenStudy (amistre64):

on which part? you are given a function and asked to define the tangent line at x=1, and use it to estimate the value of x=0.8

OpenStudy (amistre64):

the slope of a line is the derivative; and the point to anchor it to is given when x=1, then y = 1^7 = 1

OpenStudy (amistre64):

plug in 0.8 into the tangent line ... compare that to the calculator value of.8^7

OpenStudy (amistre64):

the percent of error is the difference divide by the actual value (.8)^7 - tan y ------------ (.8)^7

OpenStudy (anonymous):

urghh sorry my laptop died! why do you only plug in .8 into the tan line instead of 0.8^7?

OpenStudy (amistre64):

the tangent line is approximating the function x^7 its simpler to assess the tangent line than it is to assess the ^7 function

OpenStudy (amistre64):

the input for either one is .8 ... the tangent line is just approximating the value to a certain degree of accuracy